OpenStudy (anonymous):

What is wrong with the following calculation? int_{-2}^{1}\frac{ 1 }{ x^4 }dx=\frac{ -1 }{ 3x^3 }]^1(-2)below=\frac{ -1 }{ 3 }-\frac{ 1 }{ 24 }=\frac{ -9 }{ 24 }

OpenStudy (anonymous):

This is the question. \[\int\limits_{-2}^{1}\frac{ 1 }{ x^4 }dx\]

OpenStudy (zarkon):

what happens when x=0

OpenStudy (anonymous):

it would be undefined?

OpenStudy (zarkon):

so this integral is an improper integral

OpenStudy (anonymous):

what? I don't quite understand...

OpenStudy (zarkon):

was all that work above given and you are to tell why it is wrong?

OpenStudy (zarkon):

or is it your work?

OpenStudy (anonymous):

It's not my work, it's from the textbook

OpenStudy (zarkon):

to use the FTC the integrand (the thing inside the integral) must be continuous on the interval you are integrating over

OpenStudy (anonymous):

okay...

zepdrix (zepdrix):

still confused? :o

OpenStudy (anonymous):

yes... :(

zepdrix (zepdrix):

1/x^4 has an asymptote at x=0, ya? :o

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

so our function is not continuous over the given interval.See the problem the 0 is creating for us? +_+